UT 25 Aug 2020 - Night #4 of Restart

OSA: D.G. Huerta

ISp: E. Solheid

Mtn Manager: R. Hansen

ISA: B. Rothberg (Tucson - secret bunker)

Special Guest stars: O. Kuhn (Tucson), R. Rothberg (Tucson - secret bunker)

Problems

  1. LBC-Blue unavailable tonight - may need to remove from telescope and do swap in the lab.
  2. MODS1-Blue IC reported Seq unavailable at UT 01:08 - restarted IC via Raritan but it hung - could not run any isis commands to contact M1BC, sequenceer came back on line at UT 01:17 without my (or anyone's) intervention.
  3. DIMM still unavailable at UT 01:32
  4. Seqencer issue mods2b.2020825.0001.fits IT 6386
  5. LBC images are not going into /newdata - According to the log data we get a tftp -b32768 -o 192.168.34.15 put lbcb.20200825.025129.fits d:\lbcb.20200825.025129.fits > nul 2>&1< upload command failed with retcode:8
  6. RFBG rotator issues

Night Log:

UT 01:29 - MODS-1 and -2 awake

UT 01:50 - taking second sievesnap, mods2blue showing bad sequener on image 0001, 0002 ok

UT 03:40 - still closed for clouds - D.Thompson working on LUCI-2 functional checkout, but RFBG is having issues.

UT 03:51 - R. Rothberg arrived to check on progress

UT 04:20 - heard a boom over the polycom - sounds like thunder up there - All-sky shows a bright light.

-definitely have lightning - see it on the roofcam

UT 04:55 - troubleshooting of LBC to archive file transfer, all-sky shows rain drops, SM reports rain and 95% humidity, LBT shows 46% humidity.

UT 06:52 - going to power off and power on LBCs and try an exposure

UT 07:04 - same results

2020/08/25 07:04:35.171000 N B CAMERA CCDCTRL issuing >tftp -b32768 -o 192.168.34.15 put lbcb.20200825.045346.fits d:\lbcb.20200825.045346.fits > nul 2>&1< upload command ...
2020/08/25 07:04:35.218000 W B CAMERA CCDCTRL issuing >tftp -b32768 -o 192.168.34.15 put lbcb.20200825.045346.fits d:\lbcb.20200825.045346.fits > nul 2>&1< upload command failed with retcode:8
2020/08/25 07:04:35.233000 N B CAMERA CCDCTRL file "d:\lbcb.20200825.045346.fits" queued for upload

UT 07:05 - powering off LBCs

UT 09:57 - still closed for clouds

UT 11:15 - putting MODS to sleep

EON

PROPOSED PLAN OF ACTION:

  • MODS
    • Basic Checkout
      • Observe a standard star in all modes (30-45 min for 3 images).
        • Repeat if clear, otherwise take specstd for any modes used during the night
      • Dice-5 imaging offsets
        • Dumbell Nebula (first half)
        • M76 (2nd half)
    • Test Science OB
      • Sidereal
      • Non-Sidereal
    • Spectrophotometric standards (only if clear)
  • LBC- Red
    • DOFPIA / doHybrid - colliimatie on sky
    • Test Images
  • Mixed mode MODS-1/LBC-Red tests

Closed Dome Activities

  • MODS Prism Calibrations
    • Dual
    • Red
    • Blue
  • LUCI-2 functional check (D. Thompson)

-- BarryRothberg - 24 Aug 2020
Topic revision: r7 - 25 Aug 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback